Systems Staff Meeting Notes 7/24/00

Systems Staff Meeting Notes
July 24, 2000

Present: Cynthia Bowker, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. This is Cynthia's first staff meeting!
  2. Mystery of the Gateway RMA labels - Geoffrey is trying to return two Gateway items. The shipping labels they sent had the wrong return address - Leigh Hall instead of Smathers. He is currently in email correspondence with Gateway's Customer Support, trying to get them to send new labels with the correct address.
  3. PAULA - Try to be more specific when signing out and give an approximate time of return.
  4. Facilities Pest Reports - there is now a new procedure for reporting vermin spotted in the libraries on the Facilities Web page. Click on the service request button then log in your "pest report".
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Most of the VISORS have been distributed. There are another 15 on order, no distribution list yet.
  2. Franklin Covey company - no relation to Bill Covey. If we receive a shipment from them, it is not personal mail for Bill.
  3. HS&S laptop status - the whole unit was sent to Gateway for repair. Need to prod Gateway to return it soon since Gary will be leaving for camp and he will need it.
Liaisons
  1. The Liaisons have requested that we be more formal in scheduling repair appointments. They would like to be present for the repair as well. If it's going to take a long time, we'll need to see if they want to wait or let us proceed without them.
  2. We'll need to inform the Trouble Ticketing System of the name of the liaison that is involved on the ticket if they decide to wait.
Networking
  1. DLC - security problems again, will be changing the locks in that area. They've also had some database corruption. Erich has requested the installation of security cameras. We need to be careful when working in this area.
  2. Geoffrey is setting up micros for troubleshooting in the closets.
  3. Digital is still working in Chemistry. Bobbie talked with Judy & Tim today, they said they could be ready to come over here by next week. They are confused as to what exactly they need to be doing on the 2nd & 3rd floors - as is Bobbie. Facilities has ripped apart the columns for random furniture moving. They say they moved the columns to run the electrical. Digital will replace the copper with fiber.
  4. Closets in West - fibers were not wired to spec. We need to get a final copy of Judy's spec. We knew they were short in the pull from room 219 to 229, but there is no nine-pair in the other closets. This is wrong - we need to pull the contracts & see if we paid for nine-pairs. If the contracts were changed at the last minute, we need to find out how.
  5. Facilities says there is no money available to buy the conduit for completing the wireless in the stacks.
Software Projects Status
  1. Cynthia Bowker joins the programming staff today as a new analyst. She will be part time (afternoons) for the next four weeks, as she is still covering her old position in the Director's office.
  2. OPS tracking system mods are complete. Resource Services is in too much turmoil to do anything with them yet.
  3. Resource Services NT Chinese-Japanese-Korean connection to FCLA Z39.50 - they are making progress connecting to the server, but it still doesn't display quite right. Winston is still investigating.
  4. Dialog User Interface - Vernon demoed it, he doesn't think there is a problem. Winston will email John Freund to see if he's having trouble with it, but for now this project is on the back burner.
  5. DLC - Ying is still working on the tracking database. She should have the first phase finished by this week. She also needs to create a contact list on the SQL server to import to Outlook. May need to look at an automatic way to download to the SQL server. Need primary key to leave in Exchange - VBA may work. There's no way to hook the project to a vendor. The project could be a mailing list - global list of vendors - may dovetail with Win '2000.
  6. DLC TEI - text editing interface; not started yet. Embedded mark-up tags to make it easier to search.
Calendar
  1. Will out July 24-28.
  2. Rich returns July 26.


Back to the Systems Homepage

Last updated July 31, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu